calculate the shortest distance in which an automobile moving with a speed of 108 km/h can be stopped .If the coefficient of friction between the tires and the road is 0.74 .Also calculate the retardation .

Grade:12

1 Answers

Eshan
5 years ago
The maximum frictional force acting on a body is\mu N=\mu mg
Hence the maximum retardation of the body=\dfrac{\mu mg}{m}=\mu g=7.4m/s^2

Initial speed of the body=108km/h=30m/s

Hence the shortest distance in which the automobile stops=\dfrac{u^2}{2a}=\dfrac{30\times 30}{2\times 7.4}m=60.8m